A capillary tube with an inside radius of 638 π‘›π‘š can support a 100 π‘šπ‘š column of liquid that has a density of 820 π‘˜π‘”π‘šβˆ’3 . The observed contact angle is 12.5Β°. Find the surface tension of the liquid for β„Ž = 20 π‘π‘š.Β 


1
Expert's answer
2021-12-06T09:47:40-0500

Determine the surface tension:


"h=\frac{2\sigma\cos\theta}{\rho gr},\\\space\\
\sigma=\frac{\rho grh}{2\cos\theta},\\\space\\
\sigma=\frac{820·9.8·638·10^{-9}·0.2}{2\cos12.5}=5.25·10^{-4}\text{ N/m}."